Circular No. 6826

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998R IN UGC 6271 S. Jha, P. Garnavich and R. Kirshner, Harvard-Smithsonian Center

for Astrophysics, report the discovery by P. Berlind and B. Carter of a supernova in UGC 6271. Supernova 1998R was found on Feb. 23.5 UT with the Whipple Observatory’s 1.5-m reflector + FAST spectrograph during the course of the CfA Spectral Type Survey. A spectrum of the supernova reveals that it is of type II and after maximum: it displays a prominent P-Cyg Hα profile with an expansion velocity of 6200 km/s. The host galaxy’s recession velocity, measured from narrow emission lines, is 2006 ± 40 km/s. Confirming CCD images taken by D. Koranyi with the 1.2-m telescope yield for the supernova a precise position of α = 11^{h}14^{m}37^{s}.03, δ = +30^{o}18′51′′.7 (equinox 2000.0, uncertainty ±0′′.5), which is offset 3′′.1 west and 0′′.8 north from the center of the galaxy. The estimated apparent magnitude of the supernova is V = 16.

XTE J1858+034 R. Remillard and A. Levine, Massachusetts Institute of Technology

(MIT), report, for the RXTE ASM team at MIT and NASA Goddard Space Flight Center (GSFC): “The RXTE All Sky Monitor has detected a new, hard x-ray transient, which we designate XTE J1858+034. The source was detected in a correlation map constructed by superposing data collected during Feb. 6–12. The source is located at α = 18^{h}58^{m}41^{s}, δ = +3^{o}24′.0 (equinox 2000.0), with an estimated uncertainty (90-percent confidence) of 6′. Reanalysis of recent ASM observations shows that the source flux (2–12 keV) was 15 mCrab on Jan. 30, 24 mCrab on Feb. 7, 15 mCrab on Feb. 16 and 10 mCrab on Feb. 26. The spectrum is very hard, resembling that of an x-ray pulsar.”

T. Takeshima and R. H. D. Corbet, GSFC and Universities Space Research Association; F. E. Marshall and J. H. Swank, GSFC; and D. Chakrabarty, MIT, report: “PCA/RXTE observations of XTE J1858+034 were performed on Feb. 20 (1 ks duration) and 24 (8 ks). The source flux (2–10 keV) was 13 and 12 mCrab on Feb. 20 and Feb. 24, respectively. The identification of this source as a pulsar is confirmed by the detection of pulsations in both observations, and a period of 221.0±0.5 s was determined from the Feb. 24 observations. The pulse profile is nearly sinusoidal with a pulsed fraction of ∼ 25 percent (peak to peak). The pulsations and transient nature of XTE J1858+034 suggest that this is a Be/neutron star binary. Optical observations are strongly encouraged.”

1998 February 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Brian G. Marsden